Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, December 16, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 16, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/16/98 (Wednesday)

Segment 1 report for 12/16/98: Testing of the electrical equipment continued at Emidio Pump Station. The coating crew worked on the pipe tied-in to the Stream Crossing #4 bore from EMP 8.98 to 8.94. Pipe was being pushed through the California Aqueduct bore (EMP 12.28). At the Tecuya Creek/I-5 bore tie-in to the main line is under way from EMP 9.85 to 9.81. Casing is being installed at both the Stream Crossing #9 bore (EMP 15.35) and the Tejon Bench/I-5 bore (EMP 16.49). Shading continued on Grapevine Rd. to EMP 15.78. At Grapevine Pump Station electrical work continued. On Tejon Bench, shading continued from EMP 16.67 to 16.59. Excavation occurred to EMP 18.01. On Lebec Rd. excavation continued to EMP 21.64 (working south) and ~50 feet was excavated to EMP 21.71 (working north). Conduit was placed through the Cuddy Creek Tributary bore (EMP 25.47). Cable was pulled from EMP 20.75 to 19.70.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page